OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? My answer is 3x - y = -5 Am I right? :)

OpenStudy (anonymous):

hmm i did not get the same answer :( i got the same numbers but negative :(

OpenStudy (anonymous):

sorry was helping someone here i will do my problem step by step and you can check me so in case im wrong we know :)

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

\[\frac{1-10}{-2-1}=-3\]\[y - 1 = -3(x+2)\]\[y - 1 = -3x - 6\]\[y = -3x - 5\] check my work :)

OpenStudy (anonymous):

ops i see my mistake already :( it should be a positive 3 :O so that changes my final to y = 3x + 7

OpenStudy (anonymous):

So then it would be 3x - y = -7 ?

OpenStudy (anonymous):

yea that looks right :) srry bout my mistake !

OpenStudy (anonymous):

Thank you! :)